# Understood: Real-Time Communication Support for Adults with ADHD Using Mixed Reality

### v1

@@ 6.2. Complementing Long-Term Root-Cause Interventions with Immediate Symptomatic Support
In our user study, clinicians acknowledged the effectiveness of Understood in helping patients navigate moment-to-moment communication challenges. However, they also expressed concerns about the need to balance short-term symptomatic relief with long-term, root-cause therapeutic strategies. Understood offers immediate support mechanisms – such as conversation summarization, word suggestions, and off-topic detection – that primarily address the surface-level manifestations of ADHD. These features function like “training wheels” when learning to ride a bicycle: they compensate for current functional gaps but do not directly address the underlying cognitive causes of attention dysregulation. While it does not address the underlying causes, such symptomatic support can significantly reduce communication anxiety and social stigma, making it especially valuable for individuals who lack consistent access to structured, long-term treatment.
In contrast, interventions such as CBT and SST aim to achieve root-level change by targeting deeper cognitive mechanisms such as attention regulation, self-monitoring and strategy generation (willis2019stand; ladd1983cognitive; gokel2017effects). These therapies rely on guided feedback, structured practice, and cognitive restructuring over extended periods. Using the same metaphor, while our system helps “hold the user balanced before they fall”, traditional therapies aim to “teach the user how to maintain balance on their own”.
To strike a balance between these two modes of support, all four auxiliary functions in Understood are designed to be user-triggered, activated only when the user deliberately gazes downward at the trigger box. This design ensures that assistance is delivered only when actively sought, helping preserve cognitive engagement and user agency. Furthermore, we explicitly position Understood as a complement – not a replacement – for therapist-mediated interventions. This distinction is crucial given the phenomenon of Context Dependence (edwards2024real; el2019context), where users may perform well within structured therapy environments but struggle to apply learned strategies in unstructured, real-world scenarios. In this light, Understood may serve as a valuable daily-life support layer, bridging formal interventions and everyday application. Future work may explore how to further personalize in-situ support by aligning real-time system behavior with individualized therapeutic goals.
@@ 6.3. Generalizing Understood beyond ADHD
Our user study demonstrates that Understood effectively supports individuals with ADHD in overcoming communication difficulties, while preserving ease of use. Beyond ADHD, other neurodivergent populations – such as individuals with aphasia, autism spectrum disorder (ASD), or pragmatic language impairment (PLI) – often face analogous communicative challenges, including memory disruptions and conversational disfluency (CLARK2003265; MARKOWITSCH2003287; Wilkinson1998Profiles; Jenny2013Social). These parallels suggest the potential for Understood to be adapted to a broader range of user needs.
@@
Moreover, the applicability of Understood may extend beyond the domain of neurological disorders. For example, second-language (L2) speakers often experience similar breakdowns in fluency and conversational flow due to cognitive overload or limited linguistic resources. In such cases, Understood could act as an immersive communication scaffold – providing predictive suggestions that support word retrieval, reduce cognitive load, and foster linguistic intuition. This aligns with the principle articulated by Elise Roy in her TED talk: “When we design for disability, we all benefit” (EliseRoy).
@@ 6.4. Limitations and Future Work
Second, the performance of Understood is constrained by external technological limitations, specifically the use of the GPT-4o model. We prioritized accurate feedback over response time, which led to occasional delays in generating responses. While alternatives like GPT-3.5 Turbo (gpt3.5) and Phi-4-Mini (phi4) offer quicker response times (around 1.2 seconds and 0.6 seconds, respectively), these models may compromise on the accuracy of the feedback, producing vague or overly lengthy responses. Despite the response delays, GPT-4o provides more accurate feedback, which aligns with our system’s goals. Additionally, LLMs can still introduce hallucinations in feedback, as observed in our user study, where users noted issues with the perceived correctness of the responses. Future advancements in models like GPT o3-mini (gpto3) and DeepSeek-R1 (deepseek) may offer more accurate responses, but at the cost of longer response times, which would not meet our real-time requirements.
@@
Additionally, our study did not compare Understood with traditional interventions, such as SST and CBT, which are long-term approaches. Instead, Understood was positioned as a complementary tool for individuals who face challenges accessing these traditional therapies. Future work could explore how Understood can function as a supplement to, or even as a part of, longer-term interventions.
Third, We also acknowledge the lack of long-term evaluations. Limited by time constraints, we did not assess the users’ experiences or progress over an extended period. Future research could incorporate longitudinal studies to better understand how users’ performance and experiences evolve over time.
@@ 7. Conclusion
In this study, we introduced Understood, an MR system designed to assist adults with ADHD in real-world communication. Developed through a formative study with semi-structured interviews and a design study, Understood features three key features on (1) real-time conversation summarization, (2) context-aware subsequent word suggestions and (3) topic shifting detection and reminding. User study demonstrated that Understood effectively supports communication with high usability, offering a complement to therapist-mediated interventions. This research not only deepens our understanding of the communication challenges faced by adults with ADHD, but also prompts a dialectical reflection on the balance between external support and self-reliance, as well as the interplay between symptomatic support and root-cause interventions. We view Understood as a step towards advancing communication practices, potentially paving the way for more sophisticated human-computer collaboration in accessibility contexts.
@@ Acknowledgements.
Occurrences on Patients
@@
Table 3. The 15 statements derived from DIVA-5 regarding the communication challenges on individuals with ADHD, with the ratings by clinicians and occurrences on patients for each statement.

### v2

@@ 6.3. Complementing Long-Term Root-Cause Interventions with Immediate Symptomatic Support
In our user study, clinicians acknowledged the effectiveness of Understood in helping individuals with ADHD navigate moment-to-moment communication challenges. However, they also expressed concerns about the need to balance short-term symptomatic relief with long-term, root-cause therapeutic strategies. Understood offers immediate support mechanisms – such as conversation summarization, word suggestions, and off-topic detection – that primarily address the surface-level manifestations of ADHD. These features function like “training wheels” when learning to ride a bicycle: they compensate for current functional gaps but do not directly address the underlying cognitive causes of attention dysregulation. While it does not address the underlying causes, such symptomatic support can significantly reduce communication anxiety and social stigma, making it especially valuable for individuals who lack consistent access to structured, long-term treatment.
In contrast, interventions such as CBT and SST aim to achieve root-level change by targeting deeper cognitive mechanisms such as attention regulation, self-monitoring and strategy generation (Willis et al., 2019; Ladd and Mize, 1983; Gökel and Dağlı, 2017). These therapies rely on guided feedback, structured practice, and cognitive restructuring over extended periods. Using the same metaphor, while our system helps “hold the user balanced before they fall”, traditional therapies aim to “teach the user how to maintain balance on their own”.
To strike a balance between these two modes of support, all four auxiliary functions in Understood are designed to be user-triggered, activated only when the user deliberately gazes downward at the trigger box. This design ensures that assistance is delivered only when actively sought, helping preserve cognitive engagement and user agency. Furthermore, we explicitly position Understood as a complement – not a replacement – for therapist-mediated interventions. This distinction is crucial given the phenomenon of Context Dependence (Edwards et al., 2024; El Tohamy et al., 2019), where users may perform well within structured therapy environments but struggle to apply learned strategies in unstructured, real-world scenarios. In this light, Understood may serve as a valuable daily-life support layer, bridging formal interventions and everyday application. Future work may explore how to further personalize in-situ support by aligning real-time system behavior with individualized therapeutic goals.
@@ 6.4. Generalizing Understood beyond ADHD
Our user study demonstrates that Understood effectively supports individuals with ADHD in overcoming communication difficulties, while preserving ease of use. Beyond ADHD, other neurodivergent populations – such as individuals with aphasia, autism spectrum disorder (ASD), or pragmatic language impairment (PLI) – often face analogous communicative challenges, including memory disruptions and conversational disfluency (Clark and Cummings, 2003; Markowitsch, 2003; Wilkinson, 1998; Gibson et al., 2013). These parallels suggest the potential for Understood to be adapted to a broader range of user needs.
@@
Moreover, the applicability of Understood may extend beyond the domain of neurological disorders. For example, second-language (L2) speakers often experience similar breakdowns in fluency and conversational flow due to cognitive overload or limited linguistic resources. In such cases, Understood could act as an immersive communication scaffold – providing predictive suggestions that support word retrieval, reduce cognitive load, and foster linguistic intuition. This aligns with the principle articulated by Elise Roy in her TED talk: “When we design for disability, we all benefit” (Roy, 2016).
@@ 6.5. Limitations and Future Work
Additionally, although the HoloLens provides a comprehensive heads-up display for delivering assistive functionalities, we acknowledge that such devices tend to be highly visible, potentially disrupting the social dynamics of face-to-face interactions. This visibility may increase the risk of stigmatization for our users with ADHD, reinforcing stereotypes such as being perceived as “difficult to communicate with”. Future designs might consider alternative support platforms with lower visual intrusiveness, such as Google Glass, and could incorporate complementary modalities like subtle auditory cues to deliver assistance in a less obtrusive manner.
Third, the performance of Understood is constrained by external technological limitations, specifically the use of the GPT-4o model. We prioritized accurate feedback over response time, which led to occasional delays in generating responses. While alternatives like GPT-3.5 Turbo (Ye et al., 2023) and Phi-4-Mini (Abouelenin et al., 2025) offer quicker response times (around 1.2 seconds and 0.6 seconds, respectively), these models may compromise on the accuracy of the feedback, producing vague or overly lengthy responses. Despite the response delays, GPT-4o provides more accurate feedback, which aligns with our system’s goals. Additionally, LLMs can still introduce hallucinations in feedback, as observed in our user study, where users noted issues with the perceived correctness of the responses. Future advancements in models like GPT o3-mini (OpenAI, 2025) and DeepSeek-R1 (Guo et al., 2025) may offer more accurate responses, but at the cost of longer response times, which would not meet our real-time requirements.
@@
Additionally, our study adapted items from the NASA-TLX and related tools to assess perceived mental/physical demand, performance, and frustration, complemented by objective indicators such as response length and accuracy. These jointly support the claim that the system facilitates cognitive offloading. However, as the full NASA-TLX was not directly administered, this may limit the completeness and comparability of our assessment. We acknowledge this limitation and suggest that future work incorporate the full instrument for more standardized evaluation.
It’s also worth noting that our study did not compare Understood with traditional interventions, such as SST and CBT, which are long-term approaches. Instead, Understood was positioned as a complementary tool for individuals who face challenges accessing these traditional therapies. Future work could explore how Understood can function as a supplement to, or even as a part of, longer-term interventions.
Fourth, we also acknowledge the lack of long-term evaluations. Limited by time constraints, we did not assess the users’ experiences or progress over an extended period. Future research could incorporate longitudinal studies to better understand how users’ performance and experiences evolve over time.
@@ 7. Conclusion
In this study, we introduced Understood, an MR system designed to assist adults with ADHD in real-world communication. Developed through a formative study with semi-structured interviews and a design study, Understood features three key features on (1) real-time conversation summarization, (2) context-aware subsequent word suggestions and (3) topic shifting detection and reminding. User study demonstrated that , when evaluated under a within-subject, lab-based setting, Understood effectively supports communication with high usability, offering a complement to therapist-mediated interventions. This research not only deepens our understanding of the communication challenges faced by adults with ADHD, but also prompts a dialectical reflection on the balance between external support and self-reliance, as well as the interplay between symptomatic support and root-cause interventions. We view Understood as a step towards advancing communication practices, potentially paving the way for more sophisticated human-computer collaboration in accessibility contexts.
@@ Acknowledgements.
We thank anonymous reviewers for their valuable feedback. We would like to thank Dr. Sichu Wu at The Affiliated Brain Hospital of Nanjing Medical University, Dr. Hao Yao at Shanghai Mental Health Center and Yuchen Wu from University of Waterloo for their insightful suggestions. We would like to express our gratitude to Irasutoya [^9] https://www.irasutoya.com/ for inspiring the illustration style used in part of this paper. This work is supported by grants from the National Natural Science Foundation of China (No. 62372298), Shanghai Engineering Research Center of Intelligent Vision and Imaging, Shanghai Frontiers Science Center of Human-centered Artificial Intelligence (ShangHAI), and MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (KLIP-HuMaCo).
@@
Occurrences on Individuals with ADHD
@@
Table 3. The 15 statements derived from DIVA-5 regarding the communication challenges on individuals with ADHD, with the ratings by clinicians and occurrences on individuals with ADHD for each statement.